# Better neural network expressivity: subdividing the simplex

### v1

@@ 3.2 Upper bounds on the depth via subdivisions
$(P\cup Q)+(P\cap Q)=P+Q.$ (3)
The deeper fact we use is an inclusion-exclusion extension of the valuation property. It is called the full additivity of the identity map; see for example 5.20 in [26] and Chapter 6 in Schneider’s book [31].
@@
The proof of Lemma 9 is quite deep. First, it uses the equivalence between polytopes and their support functions. Second, it extends the support function to the lattice of finite unions of polytopes. This extension is topological and uses the Euler characteristic. Third, the Euler characteristic is shown to be fully additive. Finally, we can now deduce that the extended support function is fully additive, and hence the identity map on polytopes is fully additive (the identity map may not be fully additive on the lattice of finite unions of polytopes).
Lemma 9 is useful in combination with Lemma 8 because from a cover of $X=\Delta^{n-1}$ by polytopes $Q_{1},\ldots,Q_{m}$, we can write $\Delta^{n-1}$ as a formal Minkowski difference between two polytopes that are as complex as the $Q_{S}$’s. This is captured by the following lemma.
@@
Observe that the condition $Q_{S}\in\operatorname{\mathcal{P}}_{k}$ holds if $Q_{1},\ldots,Q_{m}\in\operatorname{\mathcal{P}}_{k}$ and each $Q_{S}$ is a face of some $Q_{i}$ because $\operatorname{\mathcal{P}}_{k}$ is closed under taking faces. This observation suggests the following way of using Lemma 10. Instead of writing $\Delta^{n-1}$ as a formal Minkowski difference, it suffices to subdivide $\Delta^{n-1}$ into $Q_{1},\ldots,Q_{m}\in\operatorname{\mathcal{P}}_{k}$. If the subdivision is polyhedral, the intersections $Q_{S}$ are always faces, and such a subdivision automatically leads to writing $\Delta^{n-1}$ as $\Delta^{n-1}+P=Q$ for $P,Q\in\operatorname{\mathcal{P}}_{k}$. Note that $P$ and $Q$ themselves might be pretty complicated and “hard to find” because they originate from the exponentially large alternating sums in Lemma 9.
3.3 Subdividing the four-simplex
We now apply our previously developed machinery to explain the geometric intuition behind our construction of the two-hidden-layer $\operatorname{\mathsf{ReLU}}$ neural network for $\mathsf{MAX}_{5}$. Geometrically, we subdivide $\Delta^{4}$ to four pieces, each in $\operatorname{\mathcal{P}}_{2}$. This subdivision leads to the two-hidden-layer $\operatorname{\mathsf{ReLU}}$ network for $\mathsf{MAX}_{5}$ described in Equation 2.
@@ 3.3 Subdividing the four-simplex
where $p_{i}\in\operatorname{\mathcal{P}}_{0}$ is a point and $Z_{i}\in\operatorname{\mathcal{P}}_{1}$ is a rhombus. In words, each $Q_{i}$ is a rhombic pyramid. This is a nice 3D puzzle, and an illustration of the solution is given in Figure 1. This subdivision leads to a computation of $\mathsf{MAX}_{4}$ with two hidden layers, which is different than the “trivial network”.
@@
The convex hull of the four points is a three-dimensional simplex (which is affinely equivalent to $\Delta^{3}$). Denote by $x_{ij}$ the midpoints of $x_{i}$ and $x_{j}$. The orthogonal projections of $x_{1},\ldots,x_{4}$ to the plane $e_{3}^{\perp}$ are the four vertices of the square $[-1,1]^{2}$. Cutting the simplex along the planes $e_{1}^{\perp}$ and $e_{2}^{\perp}$ partitions the simplex into four identical rhombic pyramids:
@@
Figure 1: Subdividing the three-simplex into four identical rhombic pyramids. The highlighted rhombic pyramid has base $x_{1},x_{14},x_{34},x_{13}$ and apex $x_{12}$.
@@
The four full-dimensional polytopes in $\operatorname{\mathcal{P}}_{2}$ in the subdivision are the Newton polytopes of the four terms $P_{1}(x)$, $P_{2}(x)$, $P_{3}(x)$, and $P_{4}(x)$ in Equation 2, while the other terms in Equation 2 are the support functions of lower-dimensional faces of the four full-dimensional polytopes.
3.4 Higher dimensions
The geometric approach also allows to see why we can compute $\mathsf{MAX}_{3^{n}+2}$ in $n+1$ layers. We have seen that $\Delta^{3}$ can be subdivided into four pieces, each of which is a join222that is, a convex hull of two polytopes that are embedded in non-intersecting affine spaces of a point $p_{i}$ and a parallelogram $Z_{i}$. The simplex $\Delta^{3}$ is the join of a point and a triangle. In other words, we replaced a triangle (which is in $\operatorname{\mathcal{P}}_{2}$) with a parallelogram (which is in $\operatorname{\mathcal{P}}_{1}$). More abstractly, we replaced three points (the triangle) by a single parallelogram, which is the geometric intuition behind 5. This suggests that each layer can reduce “the number of objects to be joined” by a factor of three, which should then allow $\Delta^{n}$ to be computed in roughly $\log_{3}(n)$ layers. A formal implementation of this proof strategy can be done, but it requires a non-trivial reasoning on the interactions between joins and subdivisions. The algebraic proof presented in Section 2.2 “hides” this complex interaction in the vector space structure of $\mathcal{T}_{a,b}$.
@@ Acknowledgments
We thank Daniel Reichman for valuable discussions around $\operatorname{\mathsf{ReLU}}$ networks. C.H. thanks Gennadiy Averkov for valuable discussions around Lemma 9 during the Cargese Workshop on Combinatorial Optimization 2024. E.B., F.B. and A.Y. are supported by a DNRF Chair grant. E.B., F.B., J.S. and A.Y. are partially supported by BARC.

### v2

@@ 3.2. Upper bounds on the depth via subdivisions
(3) $(P\cup Q)+(P\cap Q)=P+Q.$
The deeper fact we use is an inclusion-exclusion extension of the valuation property. It is called the full additivity of the identity map; see for example 5.20 in [MS83] and Chapter 6 in Schneider’s book [Sch13].
@@
It is worth mentioning that Lemma˜9 leads to a different proof of the important Equation˜1 that was proved by Wang and Sun [WS05], compare also [HBDSS23]. The explanation, in a nutshell, is as follows. Every polytope $P$ can be triangulated, that is, subdivided into simplices. Then, by the Lemma˜9, the support function $h_{P}$ can be written as a difference between the support functions of simplices. In other words, $h_{P}$ is a signed-sum of $\mathsf{MAX}_{n+1}$ functions.
The proof of Lemma˜9 is quite deep. Like the inclusion-exclusion formula, it is tempting to prove it by induction. But as opposed to the inclusion-exclusion formula, if we remove a single polytope $Q_{i}$ then the relevant union may no longer be a polytope and the induction hypothesis can not be applied. We must significantly strengthen the lemma in order to prove it. The proof is obtained in several steps. First, it uses the equivalence between polytopes and their support functions. Second, it extends the support function to the lattice of finite unions of polytopes. This extension is topological and uses the Euler characteristic. Third, the Euler characteristic is shown to be fully additive. Finally, we can now deduce that the extended support function is fully additive, and hence the identity map on polytopes is fully additive (the identity map may not be fully additive on the lattice of finite unions of polytopes).
Lemma˜9 is useful in combination with Lemma˜8 because from a cover of $X=\Delta^{n-1}$ by polytopes $Q_{1},\ldots,Q_{m}$, we can write $\Delta^{n-1}$ as a formal Minkowski difference between two polytopes that are as complex as the $Q_{S}$’s. This is captured by the following lemma.
@@
Observe that the condition $Q_{S}\in\operatorname{\mathcal{P}}_{k}$ holds if $Q_{1},\ldots,Q_{m}\in\operatorname{\mathcal{P}}_{k}$ and each $Q_{S}$ is a face of some $Q_{i}$ because $\operatorname{\mathcal{P}}_{k}$ is closed under taking faces. This observation suggests the following way of using Lemma˜10. Instead of writing $\Delta^{n-1}$ as a formal Minkowski difference, it suffices to subdivide $\Delta^{n-1}$ into $Q_{1},\ldots,Q_{m}\in\operatorname{\mathcal{P}}_{k}$. If the subdivision is polyhedral, the intersections $Q_{S}$ are always faces, and such a subdivision automatically leads to writing $\Delta^{n-1}$ as $\Delta^{n-1}+P=Q$ for $P,Q\in\operatorname{\mathcal{P}}_{k}$. Note that $P$ and $Q$ themselves might be pretty complicated and “hard to find” because they originate from the exponentially large alternating sums in Lemma˜9.
3.3. Subdividing the four-simplex
We now apply our previously developed machinery to explain the geometric intuition behind our construction of the two-hidden-layer $\operatorname{\mathsf{ReLU}}$ neural network for $\mathsf{MAX}_{5}$. Geometrically, we subdivide $\Delta^{4}$ to four pieces, each in $\operatorname{\mathcal{P}}_{2}$. This subdivision leads to the two-hidden-layer $\operatorname{\mathsf{ReLU}}$ network for $\mathsf{MAX}_{5}$ described in Equation˜2.
@@ 3.3. Subdividing the four-simplex
where $p_{i}\in\operatorname{\mathcal{P}}_{0}$ is a point and $Z_{i}\in\operatorname{\mathcal{P}}_{1}$ is a rhombus. In words, each $Q_{i}$ is a rhombic pyramid. This is a nice 3D puzzle, and an illustration of the solution is given in Figure˜1 and Figure˜2. This subdivision leads to a computation of $\mathsf{MAX}_{4}$ with two hidden layers, which is different than the “trivial network”.
@@
The convex hull of the four points is a three-dimensional simplex (which is affinely equivalent to $\Delta^{3}$). Denote by $x_{ij}$ the midpoints of $x_{i}$ and $x_{j}$. The orthogonal projections of $x_{1},\ldots,x_{4}$ to the plane $e_{3}^{\perp}$ are the four vertices of the square $1,1]^{2}$. Cutting the simplex along the planes $e_{1}^{\perp}$ and $e_{2}^{\perp}$ partitions the simplex into four identical rhombic pyramids:
@@
Figure 1. Two ways to look at the three-simplex. The highlighted part is one of four pyramids of subdivision — a rhombic pyramid with base $x_{1},x_{14},x_{34},x_{13}$ and apex $x_{12}$.
Figure 2. Subdividing the three-simplex into four identical rhombic pyramids.
@@
The four full-dimensional polytopes in $\operatorname{\mathcal{P}}_{2}$ in the subdivision are the Newton polytopes of the four terms $P_{1}(x)$, $P_{2}(x)$, $P_{3}(x)$, and $P_{4}(x)$ in Equation˜2, while the other terms in Equation˜2 are the support functions of lower-dimensional faces of the four full-dimensional polytopes.
3.4. Higher dimensions
The geometric approach also allows to see why we can compute $\mathsf{MAX}_{3^{n}+2}$ in $n+1$ layers. We have seen that $\Delta^{3}$ can be subdivided into four pieces, each of which is a join [^2] A convex hull of two polytopes that are embedded in non-intersecting affine spaces. of a point $p_{i}$ and a parallelogram $Z_{i}$. The simplex $\Delta^{3}$ is the join of a point and a triangle. In other words, we replaced a triangle (which is in $\operatorname{\mathcal{P}}_{2}$) with a parallelogram (which is in $\operatorname{\mathcal{P}}_{1}$). More abstractly, we replaced three points (the triangle) by a single parallelogram, which is the geometric intuition behind ˜5. This suggests that each layer can reduce “the number of objects to be joined” by a factor of three, which should then allow $\Delta^{n}$ to be computed in roughly $\log_{3}(n)$ layers. A formal implementation of this proof strategy can be done, but it requires a non-trivial reasoning on the interactions between joins and subdivisions. The algebraic proof presented in Section˜2.2 “hides” this complex interaction in the vector space structure of $\mathcal{T}_{a,b}$.
4. Conclusions
Summarizing, we provide the first improvement of the upper bound on the required depth to represent CPWL functions, after the initial representation result by Arora, Basu, Mianjy, and Mukherjee [ABMM18]. Thereby, we break the $\log_{2}(n)$-barrier that was conjectured to be optimal [HBDSS23] and proven to be tight in many special cases [HHL23, Val24, GHL25]. Consequently, our result has the potential to shift the focus of future work from lower bounds to upper bounds, from impossibility results to constructions. One interpretation of our result is that neural networks can compute exact maxima more efficiently than in a pairwise manner. Besides the result itself, a key contribution of our paper is that new upper bounds can be proven via subdividing the simplex into polytopes with known depth complexity, which is captured by Lemma˜10.
Curiously, while we were able to give a positive answer to the previously open question of whether the maximum of five numbers can be computed with two hidden layers, the same question remains open for the maximum of six numbers. The lower bound by Averkov, Hojny, and Merkert [AHM25] for neural networks with rational weights, however, tells us that further improving the upper bound would require rational weights with large denominators (or even irrational weights) in the neural network. Geometrically, when applying our subdivision approach, this would correspond to subdivisions of the simplex into even smaller cells. Overall, the question of whether constant depth, perhaps even two hidden layers, suffices to represent all CPWL functions remains an intriguing mystery.
@@ Acknowledgments
We thank Daniel Reichman for valuable discussions around $\operatorname{\mathsf{ReLU}}$ networks. C.H. thanks Gennadiy Averkov for valuable discussions around Lemma˜9 during the Cargese Workshop on Combinatorial Optimization 2024. E.B., F.B., and A.Y. are supported by a DNRF Chair grant. E.B., F.B., J.S., and A.Y. are part of BARC, Basic Algorithms Research Copenhagen, supported by the VILLUM Foundation grant 54451.